OpenStudy (anonymous):

Hank bought 8 hot dogs for $1.11 each and 4 drinks for $1.36 each. How much was his total purchase before tax?

OpenStudy (anonymous):

(8x1.11) + (4x 1.36) = 8.88+ 5.44 = 14.32